# Understanding uncertainty in Bayesian cluster analysis

### v1

@@ 2.2 Inference
To provide a single, representative clustering solution, the optimal Bayes estimator minimizes the posterior expected loss. For a partition-specific loss, Wade and Ghahramani (2018) recommend the Variation of Information (VI) Meilă (2007), which defines the distance between any two partitions $\rho_{1}$ and $\rho_{2}$ as:
@@
$\rho^{*}=\min_{\rho\in\mathcal{P}}\mathbb{E}\left[\text{VI}(\rho,\rho^{\prime})\mid\bm{y}\right]\approx\min_{\rho\in\mathcal{P}}\frac{1}{T}\sum_{t=1}^{T}\text{VI}(\rho,\rho^{(t)}).$
While direct minimization cannot be performed, effective algorithms exist Wade and Ghahramani (2018); Rastelli and Friel (2018); Dahl et al. (2022).
@@ 3 Methods
Summarizing the posterior with a single optimal partition not only disregards uncertainty in the clustering structure, but can also be misleading in the case of multimodality (see Example 2). To address these issues, we propose to summarize with not one, but multiple optimal partitions. Our approach borrows ideas from Balocchi et al. (2023) to approximate the posterior $\pi(\rho|\bm{y})$ with a simpler discrete distribution $q=\sum_{\ell=1}^{L}w_{\ell}\delta_{\rho_{\ell}}$ supported on a small number $L$ of partitions. Balocchi et al. focus on minimizing the Kullback-Leibler (KL) divergence and show that the optimal distribution $q$ is supported on the set of $L$ partitions that have largest posterior mass (for $L=1$, this reduces to the MAP partition). However, such partitions tend to be very similar to each other, with only one or two points having a different cluster allocation. For example, in Figure 1(a), even with the choice of $L=501$, the support of $q$ consists of the partition with a single large cluster of $500$ points and all partitions with one large cluster of size $499$ and one singleton. This would only summarize uncertainty around the mode, not the variability in the posterior.
Instead, we replace the KL divergence with the Wasserstein distance, which allows the use of metrics on the space of partitions. Here, we focus on the VI metric, but this can be extended to other choices. To formally define the Wasserstein-VI distance, let $(\mathcal{P},d_{\text{VI}})$ be the metric space on the set of partitions $\mathcal{P}$ embedded with the VI metric $d_{\text{VI}}$. Let $p$ and $q$ represent two distributions on the space of partitions $(\mathcal{P},d_{\text{VI}})$. If $\mathcal{J}(p,q)$ is the set of couplings of $p$ and $q$, i.e. the collection of distributions $J(\rho,\rho^{\prime})$ on $\mathcal{P}\times\mathcal{P}$ with marginals $p$ and $q$ on the first and second factor respectively, then we define the Wasserstein distance as
@@
In the general case of $L>1$, we can rewrite the problem of finding the WASABI posterior using properties of the Wasserstein distance between two discrete distributions.
@@
where $\mathcal{N}_{\ell}=\{\rho:d_{\text{VI}}(\rho,\rho^{*}_{\ell})<d_{\text{VI}}(\rho,\rho^{*}_{\ell^{\prime}})\textrm{ for all }\ell^{\prime}\neq\ell\}$ corresponds to the set of partitions that are closer to $\rho^{*}_{\ell}$ than to any other center. We shall name $\mathcal{N}_{\ell}$ the “region of attraction” of the center $\rho^{*}_{\ell}$. The optimal weights $w^{*}_{\ell}$ associated to each center are found by $w^{*}_{\ell}=\sum_{\rho\in\mathcal{N}_{\ell}}\pi(\rho|\bm{y}).$
In other words, the WASABI posterior is supported on the centers $\rho^{*}_{\ell}$ that minimize (3), and the probability $w^{*}_{\ell}$ associated with each of them is equal to the posterior mass of its region of attraction. See Appendix A for a proof of Proposition 1. Note that this result is known as a general property of the Wasserstein distance, see for example Section 3.1 of Cuturi and Doucet (2014).
In practice, even for moderate $n$, the objective function in (3) is intractable due to the dimension of $\mathcal{P}$. A similar approximation of the posterior can however be considered when an estimate of the posterior distribution is provided through MCMC samples. Let $\hat{\pi}(\rho|\bm{y})$ be the approximation to the posterior provided by the MCMC draws, $\hat{\pi}(\rho|\bm{y})=\frac{1}{T}\sum_{t=1}^{T}\delta_{\rho^{(t)}}$. We can then approximate $\hat{\pi}(\rho|\bm{y})$ with $q$ in a Wasserstein-VI sense.
@@
Given a set $\{\rho^{(1)},\ldots,\rho^{(T)}\}$ of MCMC draws from the posterior $\pi(\cdot|\bm{y})$, the WASABI posterior is the discrete distribution $q^{*}=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}\in\mathcal{Q}_{L}$ that is closest in a Wasserstein-VI sense to the empirical approximation of the posterior, $\hat{\pi}(\rho|\bm{y})=\frac{1}{T}\sum_{t=1}^{T}\delta_{\rho^{(t)}}$,
@@
where $\widehat{\mathcal{N}}_{\ell}=\{\rho^{(t)}:d_{\text{VI}}(\rho^{(t)},\rho^{*}_{\ell})<d_{\text{VI}}(\rho^{(t)},\rho^{*}_{\ell^{\prime}})\textrm{ for all }\ell^{\prime}\neq\ell\}$; we refer to $\widehat{\mathcal{N}}_{\ell}$ as the “empirical region of attraction” of $\rho^{*}_{\ell}$. The optimal weights $w^{*}_{\ell}$ associated to each center are proportional to the size of $\widehat{\mathcal{N}}_{\ell}$, i.e. $w^{*}_{\ell}=|\widehat{\mathcal{N}}_{\ell}|$ / T.
As described by Canas and Rosasco (2012) and first suggested by Pollard (1982), an interesting connection arises between the k-means algorithm and the approximation in a Wasserstein sense (equipped with the Euclidean distance) of an empirical distribution with a discrete distribution supported on $k$ points. In particular, the k-means algorithm provides a locally optimal solution to this problem. In our framework, we can obtain a locally optimal set of partitions that minimize (5) by applying a k-medoids-like algorithm, which generalizes k-means beyond the Euclidean distance, to the set of MCMC samples $\{\rho^{(t)}:t=1,\ldots,T\}$, with $k=L$ and distance equal to the VI distance. In other words, the MCMC-sampled partitions are clustered into $L$ groups, and each of these groups is represented by one of the centers $\rho^{*}_{\ell}$. These same centers are the support of the WASABI posterior $q^{*}$, with weights determined by the size of each group. While clustering a set of partitions may seem tautological, the value of this approach lies in its ability to summarize a large set of discrete objects with a smaller (but greater than one) number of points, each of which represents a different group or “cluster” of partitions.
@@
Consider the slightly bimodal example from the Introduction (Figure 1), with data generated from an equal-weight mixture of two unit-variance Gaussians (means $+1.1$ and $-1.1$). The minVI estimator identifies only one cluster, yet the posterior similarity matrix (PSM) suggests a possible two-cluster structure. Approximating the posterior with WASABI (using two particles), we find one particle corresponding to the single-cluster partition ($w_{1}\approx 0.7$), and another recovers the two similarly-sized clusters. In this sense, WASABI is able to recover the bimodal structure of the posterior, visible in Figure 1(a).
@@ 3.1 Algorithm
The search algorithm to find the $L$ centers that support the approximate WASABI posterior has a similar structure to the k-medoids algorithm. Essentially, given an initialization of the centers, the algorithm alternates between what we refer to as the N-update step, where the regions of attraction are updated, and the VI-search step, where the new center for each region of attraction corresponds to the minVI of each group.
More precisely, the N-update step assigns all points (the MCMC samples) to their closest center; this requires computing the VI distance between each point and each center. In the rare case when a partition is equally distant to two or more centers, it is randomly allocated among them. It is then followed by an optional outlier-search step, which is attempted when one region of attraction contains only one partition, meaning that the current center is not the closest to any other partition; this step tries to swap the current center with a sampled “outlier” partition, i.e. a partition that is far (in VI sense) from all centers, and checks if the objective function improves. Computing the VI distance between all points and centers is the most expensive part of this step, with complexity $O(nTL)$.
Next, the VI-search updates the center of each group given their new region of attraction, but unlike k-medoids, which restricts the search to the points within the group, we make use of more advanced algorithms to minimize the posterior expected VI, such as SALSO (Dahl et al., 2022). The complexity of this step depends on the algorithm selected, e.g. for SALSO (the default), the cost is $O(nTK_{\text{max}}^{2})$, where $K_{\text{max}}$ is the maximum number of clusters in the MCMC samples.
The algorithm iterates between these steps until a maximum number of iterations is reached, or the improvement in the Wasserstein distance is less than a given threshold $\epsilon$. Note that, since the VI distance is bounded above by $\log_{2}(n)$, we recommend choosing a threshold that depends on this quantity, such as $\epsilon=0.0001*\log_{2}(n)$. The basic algorithm is outlined in Algorithm 1, and a more detailed version can be found in Section B of the Appendix. As the algorithm only guarantees convergence to a local optimum, it is recommended to run it multiple times with different starting conditions. We consider several initialization strategies, including using the top $L$ partitions found with hierarchical clustering based on the PSM, or a strategy mimicking the k-means++ initialization (Vassilvitskii and Arthur, 2006), which randomly initializes the centers by selecting more spread out partitions among the samples. Further initialization details are provided in the Appendix (Section B). To reduce computational complexity, strategies employed in k-means can be adapted; for example, we allow a “mini-batch” option, where at each iteration, we consider a subsample of the MCMC partitions of size $B$. This reduces the cost to $O(nB(L+K_{\text{max}}^{2}))$ but is followed by a small number of iterations using the full set of MCMC partitions, similar to the two-stage k-means of Salman et al. (2011).
@@
To approximate the posterior distribution with WASABI, the practitioner must select the number of particles $L$. As $L\rightarrow T$, the Wasserstein distance reduces to zero, but it becomes more and more difficult to describe all centers. To achieve a balance between parsimony (small $L$) and a good enough approximation (small Wasserstein distance), we propose running the algorithm for a range of values of $L$ (usually from 1 to 10), and plotting the Wasserstein distance as a function of $L$. This produces an “elbow plot”, and the selected value of $L$ is the smallest that corresponds to a significant reduction of the Wasserstein distance, before the curve flattens (or decreases less substantially). See e.g. the left panel in Figure 2(c).
@@ 3.2 Describing and visualizing the particles
Clustering is often used for exploratory data analysis, and visualization of the partitions is a crucial part of the analysis to uncover and better understand patterns in the data. Similarly, visualization of the obtained particles is essential to better understand the structure of the posterior distribution.
For the simple slightly-bimodal data (Example 1), comparison of the particles is quite straightforward. We now consider a two-dimensional extension of this simple example, and in the following, outline several techniques to describe and visualize the particles that summarize the posterior, illustrated on this two-dimensional example.
@@
Consider a two-dimensional Gaussian mixture with four components, each with mean (±$(\pm m,\pm m)$ located in one of the four quadrants and diagonal covariance matrix with unit variance. Figure 2(a) displays the data colored by the partition used in the data-generating process (for $m=1.3$). MCMC samples are obtained by fitting a diagonal location-scale Dirichlet process mixture (DPM) using the BNPmix R package (Corradin et al., 2021). The minVI partition (Figure 2(b)) merges three of the four components, which does not reflect the axis-aligned elliptical clusters of the DPM model. This highlights that a single-point estimate can be misleading when the posterior is multimodal. In fact, the WASABI elbow plot (Figure 2(c)) suggests that multiple estimators are useful for describing the posterior. We now provide several approaches for visualizing the WASABI particles.
@@
Determining the number of clusters is an important problem, and with WASABI, we can compare the number of clusters identified by the different particles. In several experiments, we have found that different particles often tend to have different numbers of clusters. Moreover, the weights associated with the particles convey their importance in summarizing the posterior MCMC samples, as they correspond to the proportion of samples that are closer to each given particle. If a particle has a weight of 0.7, it means that 70% of the MCMC samples are closer to this particle, and can then be summarized by it. See the left panel of Figure 1(d) for a representation of these quantities for Example 1.
@@
(a) Particles’ cluster assignment.
@@
Figure 3: Several possible visualizations of the WASABI summary for Example 2: (a) particle’s scatterplot with cluster assignment, (b) posterior similarity matrix for the samples within each region of attraction, (c) the particle’s meet, (d) posterior similarity matrix using the WASABI approximation collapsed to meet’s clusters, (e) comparison of two particles using VI contribution.
Comparing pairs of particles: visualizing the VI-contribution
@@
Coloring the data points based on their VI-contribution can help pinpoint areas within the dataset where the two particles differ, enabling more detailed investigation. Figure 3(e) compares particles 2 and 3, with the lowest VI contribution for the bottom right quadrant, highlighting that these points are mostly clustered together, and the highest for points on the border, suggesting more drastic changes in who they are clustered with. See also the Appendix, for an alternative visualization.
Since the VI contribution is constant for all points that belong to the same meet cluster, we can define the “VI contribution by group” (VICG), i.e. the aggregate VIC for all points belonging to the same meet cluster.
@@
The “groups” of the VI contribution by group correspond to the meet’s clusters, which are also found as the pairwise intersection of each of the particles’ clusters. While inspecting the VIC for each data point highlights the subtle differences between two partitions, by plotting the VICG we can easily highlight high-level changes in the two partitions. In fact, major changes in the cluster assignment, such as a large cluster being split into two equally-sized subclusters, would have a significant contribution to the VI, but if the number of points in those clusters is large, the individual contribution may be negligible, since each point is still clustered together with many of the same points. However, by aggregating the contribution across all points in the meet’s cluster in the VICG, these large-scale differences become more visible. Figure 10 in the Appendix displays the VICG for the data in Example 2.
@@
Determining the uncertainty in each region of attraction: inspecting each particle’s posterior similarity matrices
The WASABI particles summarizes the posterior distribution with several “point estimates” that can be visualized for easy understanding of the posterior structure. Uncertainty remains around each of the particles, as they are the representative partition for a group of MCMC samples (their “region of attraction”). To understand the uncertainty in each region of attraction, we can visualize the PSM of each region, i.e. calculating it only using the MCMC samples that are closest to the corresponding particle. For example, Figure 3(b) displays the PSM for each of the particles’ regions of attraction for Example 2, using heatmaps by grouping (and labeling) each point into their cluster assignment. Additionally, for each of the particles we can compute the (normalized) Expected VI distance, where the expectation is taken only with respect to the MCMC samples in their region of attraction. This quantifies the spread of the samples around the particle; for example, there is more variability in the first region of attraction (characterized by single, large cluster) compared to the second (see plot titles in Figure 3(b)). An alternative visualization that groups data points by the meet’s cluster allocations is provided in Figure 8 of the Appendix.
Determining the uncertainty in each point’s cluster allocation: the contribution to the expected-VI
@@
$\text{EVI}(\rho^{*})=\sum_{i=1}^{n}\text{EVIC}_{i}(\rho^{*}).$
Note that computing the contribution of the $i$th data point the EVI of $\rho^{*}$ requires computing the expectation with respect to the posterior over the space of partitions. This can be approximated either based on the MCMC draws:
or based on the WASABI posterior:
Approximating the contribution of the $i$th data point to the EVI based on the WASABI posterior has a clear computational advantage ($O(nL)$ as opposed to $O(nT)$ with $L\ll T$); in addition, note that when $\rho^{*}$ is one of the particles, every data point in the same cluster of the meet will have the same contribution to the VI, further reducing computations and enhancing interpretation. For example, clusters in the meet with a WASABI approximated EVI of zero, are always clustered together and never with any other data points, across all particles. Lastly, we note that the EVI of the $l$th particle, $\rho^{*}_{l}$ can also be taken with respect to the posterior restricted to its region of attraction to shed light on which points are more uncertain (or stable) in their cluster allocation within the region of attraction.
@@
The contribution to the expected VI for the meet of the particles is displayed in Figure 3(c) (bottom). By analyzing the patterns in this figure, we notice that the points with the largest EVIC are the ones at the border between the four large clusters (depicted in Figure 3(c) (top) and represented here with different shapes). Moreover, the points in the top-left component (labeled with 1) have a slightly larger EVIC than the points in the bottom-right component (labeled with 6), reflecting that their cluster allocation changes more across the MCMC samples, while the cluster assignment of the points in cluster 6 change less.

### v2

@@ 2.2 Inference
To provide a single, representative clustering solution, the optimal Bayes estimator minimizes the posterior expected loss. For a partition-specific loss, Wade and Ghahramani (2018) recommend the Variation of Information (VI) (Meilă, 2007), which for any two partitions $\rho_{1}$ and $\rho_{2}$ is defined as:
$d_{\text{VI}}(\rho_{1},\rho_{2})$ $=\text{H}(\rho_{1})+\text{H}(\rho_{2})-2\text{MI}(\rho_{1},\rho_{2})$
@@
$\rho^{*}=\min_{\rho\in\mathcal{P}}\mathbb{E}\left[d_{\text{VI}}(\rho,\rho^{\prime})\mid\bm{y}\right]\approx\min_{\rho\in\mathcal{P}}\frac{1}{T}\sum_{t=1}^{T}d_{\text{VI}}(\rho,\rho^{(t)}).$
While direct minimization cannot be performed, effective algorithms exist (Wade and Ghahramani, 2018; Rastelli and Friel, 2018; Dahl et al., 2022).
@@ 3 Methods
Summarizing the posterior with a single partition not only disregards uncertainty in the clustering structure, but can also be misleading in the case of multimodality (see Example 2). To address these issues, we propose to summarize with not one, but multiple optimal partitions. Our approach borrows ideas from Balocchi et al. (2023) to approximate the posterior $\pi(\rho|\bm{y})$ with a simpler discrete distribution $q=\sum_{\ell=1}^{L}w_{\ell}\delta_{\rho_{\ell}}$ supported on a small number $L$ of partitions. Balocchi et al. focus on minimizing the Kullback-Leibler (KL) divergence and show that the optimal distribution $q$ is supported on the set of $L$ partitions that have largest posterior mass (for $L=1$, this reduces to the MAP partition). However, such partitions tend to be very similar to each other, with only one or two points having a different cluster allocation. For example, in Figure 1(b), even with the choice of $L=501$, the support of $q$ consists of the partition with a single large cluster of $500$ points and all partitions with one large cluster of size $499$ and one singleton. This would only summarize uncertainty around the mode, not the variability in the posterior.
Instead, we replace the KL divergence with the Wasserstein distance, which allows the use of metrics on the space of partitions. Here, we focus on the VI metric, but this can be extended to other choices. To formally define the Wasserstein-VI distance, let $(\mathcal{P},d_{\text{VI}})$ be the metric space on the set of partitions $\mathcal{P}$ embedded with the VI metric $d_{\text{VI}}$. Let $p$ and $q$ represent two distributions on $\mathcal{P}$. If $\mathcal{J}(p,q)$ is the set of couplings of $p$ and $q$, i.e. the collection of distributions $J(\rho,\rho^{\prime})$ on $\mathcal{P}\times\mathcal{P}$ with marginals $p$ and $q$ on the first and second factor respectively, then the Wasserstein distance is
@@
In the general case of $L>1$, we can rewrite the problem of finding the WASABI posterior using properties of the Wasserstein distance between two discrete distributions (Cuturi and Doucet (2014)).
@@
[…]
In other words, the WASABI posterior is supported on the centers $\rho^{*}_{\ell}$ that minimize (3), and the probability $w^{*}_{\ell}$ associated with each of them is equal to the posterior mass of its region of attraction. See Appendix A for a proof of Proposition 1. In practice, even for moderate $n$, the objective function in (3) is intractable due to the dimension of $\mathcal{P}$. A similar approximation of the posterior can however be considered when an estimate of the posterior distribution is provided through MCMC samples. Let $\hat{\pi}(\rho|\bm{y})$ be the approximation to the posterior provided by the MCMC draws, $\hat{\pi}(\rho|\bm{y})=\frac{1}{T}\sum_{t=1}^{T}\delta_{\rho^{(t)}}$. We can then approximate $\hat{\pi}(\rho|\bm{y})$ with $q$ in a Wasserstein-VI sense.
@@
Given a set $\{\rho^{(1)},\ldots,\rho^{(T)}\}$ of draws from the posterior $\pi(\cdot|\bm{y})$, the WASABI posterior is the discrete distribution $q^{*}=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}\in\mathcal{Q}_{L}$ that is closest in a Wasserstein-VI sense to the empirical approximation of the posterior, $\hat{\pi}(\rho|\bm{y})=\frac{1}{T}\sum_{t=1}^{T}\delta_{\rho^{(t)}}$,
@@
[…]
As described by Canas and Rosasco (2012) and first suggested by Pollard (1982), an interesting connection arises between the k-means algorithm and the approximation in a Wasserstein sense (equipped with the Euclidean distance) of an empirical distribution with a discrete distribution supported on $k$ points. In particular, the k-means algorithm provides a locally optimal solution to this problem. In our framework, we can obtain a locally optimal set of partitions that minimize (5) by applying a k-medoids-like algorithm, which generalizes k-means beyond the Euclidean distance, to the set of MCMC samples $\{\rho^{(t)}:t=1,\ldots,T\}$, with $k=L$ and distance equal to the VI. In other words, the MCMC-sampled partitions are clustered into $L$ groups, and each of these groups is represented by one of the centers $\rho^{*}_{\ell}$. These same centers are the support of the WASABI posterior $q^{*}$, with weights determined by the size of each group. While clustering a set of partitions may seem tautological, the value lies in its ability to summarize a large set of discrete objects with a much smaller number of points, each of which represents a different group or “cluster” of partitions.
@@
Consider the slightly bimodal example in Figure 1, with data generated from an equal-weight mixture of two unit-variance Gaussians (means $\pm 1.1$). The minVI estimator identifies only one cluster, yet the posterior similarity matrix (PSM) suggests a possible two-cluster structure. Approximating the posterior with WASABI (using two particles), we find one particle corresponding to the single-cluster partition ($w_{1}\approx 0.503$) and another recovers two similarly-sized clusters. In this sense, WASABI is able to recover the bimodal structure of the posterior, visible in Figure 1(b).
@@ 3.1 Algorithm
The search algorithm to find the $L$ centers that support the WASABI posterior has a similar structure to the k-medoids algorithm. Essentially, given an initialization of the centers, the algorithm alternates between what we refer to as the N-update step, where the regions of attraction are updated, and the VI-search step, where the new center for each region of attraction corresponds to the minVI of each group.
More precisely, the N-update step assigns each point (MCMC sample) to its closest center by computing the VI distance between every point and each center. In the rare case when a partition is equally distant to multiple centers, it is randomly assigned. Computing VI distances between all points and centers has complexity $O(nTL)$.
Next, the VI-search updates the center of each group given their new region of attraction, but unlike k-medoids, which restricts the search to the points within the group, we make use of more advanced algorithms to minimize the posterior expected VI, such as SALSO (Dahl et al., 2022). The complexity of this step depends on the algorithm selected, e.g. for SALSO (the default), the cost is $O(nTK_{\text{max}}^{2})$, where $K_{\text{max}}$ is the maximum number of clusters considered in the search algorithm.
The algorithm iterates between these steps until a maximum number of iterations is reached, or the improvement in the Wasserstein distance is less than a given threshold $\epsilon$. Since the VI distance is bounded above by $\log_{2}(n)$, we recommend choosing a threshold based on this, such as $\epsilon=0.0001*\log_{2}(n)$. The basic algorithm is outlined in Algorithm1, and a more detailed version is in SectionB of the Appendix. As the algorithm only guarantees convergence to a local optimum, we recommend running it multiple times with different starting conditions. We consider several initialization strategies, including using the top $L$ partitions found with hierarchical clustering based on the PSM, or a strategy mimicking k-means++ initialization (Vassilvitskii and Arthur, 2006), which randomly initializes centers by selecting more spread out partitions among the samples. (Appendix, SectionB). To reduce computational complexity, strategies used in k-means can be adapted; for example, we allow a “mini-batch” option, where at each iteration, we consider a subsample of the MCMC partitions of size $B$. This reduces the cost to $O(nB(L+K_{\text{max}}^{2}))$ but is followed by a small number of iterations using the full set of MCMC partitions, similar to the two-stage k-means of Salman et al. (2011).
@@
To approximate the posterior distribution with WASABI, the practitioner must select the number of particles $L$. As $L\rightarrow T$, the Wasserstein distance reduces to zero, but it becomes more difficult to describe all centers. To achieve a balance between parsimony (small $L$) and a good enough approximation (small Wasserstein distance), we propose running the algorithm for a range of values of $L$ (usually from 1 to 10), and plotting the Wasserstein distance as a function of $L$. This produces an “elbow plot”, and the selected value of $L$ is the smallest that corresponds to a significant reduction of the Wasserstein distance, before the curve flattens (or decreases less substantially). See e.g. Figure 2(c).
@@ 3.2 Describing and visualizing the particles
Clustering is often used for exploratory data analysis, and visualization of the partitions is a crucial part of the analysis to uncover patterns in the data. Similarly, visualization of the obtained particles is essential to better understand the structure of the posterior distribution. For the slightly bimodal data (Example 1), comparison of the particles is straightforward. We now consider a two-dimensional extension of this simple example, and in the following, outline several techniques to describe and visualize the particles that summarize the posterior, illustrated on this two-dimensional example.
@@
Consider a two-dimensional Gaussian mixture with four components, each with mean $(\pm m,\pm m)$ located in one of the four quadrants and diagonal covariance matrix with unit variance. Figure 2(a) displays the data colored by the partition used in the data-generating process (for $m=1.25$). MCMC samples are obtained by fitting a diagonal location-scale Dirichlet process mixture (DPM) using the BNPmix R package (Corradin et al., 2021). The minVI partition (Figure 2(b)) merges three of the four components, which does not reflect the axis-aligned elliptical clusters of the DPM model. This highlights that a single-point estimate can be misleading when the posterior is multimodal. In fact, the WASABI elbow plot (Figure 2(c)) suggests that multiple estimators are useful for describing the posterior. We now provide several approaches for visualizing the WASABI particles.
@@
Determining the number of clusters is an important problem, and with WASABI, we can compare the number of clusters identified by the different particles. In several experiments, we have found that different particles often tend to have different numbers of clusters. Moreover, the weights associated with the particles convey their importance in summarizing the posterior MCMC samples, as they correspond to the proportion of samples that are closer to each given particle. See the left panel of Figure 1(d) for a representation of these quantities for Example 1.
@@
(a) Data colored by particles’ cluster assignment.
@@
Figure 3: Several possible visualizations of the WASABI summary for Example 2: (a) scatterplot of the data colored by particles’ cluster assignment, (b) posterior similarity matrix for the samples within each region of attraction, (c) the particles’ meet, (d) posterior similarity matrix using the WASABI approximation collapsed to meet’s clusters, (e) comparison of two particles using VI contribution.
Comparing pairs of particles: the VI-contribution
@@
Coloring the data points based on their VI-contribution can help pinpoint areas within the dataset where the two particles differ, enabling more detailed investigation. Figure 3(e) compares particles 2 and 3, with the lowest VI contribution for the bottom right quadrant, highlighting that these points are mostly clustered together, and the highest for points on the border, suggesting more drastic changes in who they are clustered with.
Since the VI contribution is constant for all points in the same meet cluster, we can define the “VI contribution by group” (VICG), i.e. the aggregate VIC for all points in the same meet cluster.
@@
The “groups” of the VI contribution by group correspond to the meet’s clusters, which are also found as the pairwise intersection of each of the particles’ clusters. While inspecting the VIC for each data point highlights the subtle differences between two partitions, plotting the VICG highlights high-level changes in the two partitions. In fact, major changes in the cluster assignment, such as a large cluster being split into two equally-sized subclusters, would have a significant contribution to the VI, but if the number of points in those clusters is large, the individual contribution may be negligible, since each point is still clustered together with many of the same points. However, by aggregating the contribution across all points in the meet’s cluster in the VICG, these large-scale differences become more visible. Figure 11 in the Appendix displays the VICG for Example 2.
@@
Determining uncertainty in each region of attraction: inspecting each particle’s PSM
The WASABI particles summarize the posterior with several “point estimates” that can be visualized for easy understanding of the posterior structure. Uncertainty remains around each of the particles, as they are the representative partition for a group of MCMC samples (their “region of attraction”). To understand the uncertainty in each region of attraction, we can visualize the PSM of each region, i.e. calculating it only using the MCMC samples that are closest to the corresponding particle. For example, Figure 3(b) displays the PSM for each particle’s region of attraction for Example 2, using heatmaps that group points by their cluster assignment. Additionally, for each particle, we compute the (normalized) expected VI restricted to its region of attraction, i.e. approximated using only MCMC samples in its region of attraction. This quantifies the spread of the samples around the particle; for example, there is more variability in the first region of attraction (characterized by a single, large cluster) compared to the second (see plot titles in Figure 3(b)). For an alternative visualization that groups data points by the meet clusters, see Figure 9 of the Appendix.
Determining uncertainty in each point’s cluster allocation: the expected VI contribution
@@
where the expectation is taken with respect to $\pi(\rho\mid\bm{y})$, the posterior distribution over the space of clusterings. The EVI of $\rho^{*}$ is $\text{EVI}(\rho^{*})=\sum_{i=1}^{n}\text{EVIC}_{i}(\rho^{*}).$
Note that computing the contribution of the $i$th data point the EVI of $\rho^{*}$ requires computing the expectation with respect to the posterior over the space of partitions. This can be approximated either based on the MCMC draws or the WASABI posterior (Section C.1 of the Appendix). The latter has a clear computational advantage ($O(nL)$ as opposed to $O(nT)$ with $L\ll T$); in addition, note that when $\rho^{*}$ is one of the particles, every data point in the same cluster of the meet will have the same contribution to the EVI, further reducing computations and enhancing interpretation. For example, clusters in the meet with a WASABI approximated EVI of zero, are always clustered together and never with any other data points, across all particles. Lastly, we note that the EVI of the $l$th particle, $\rho^{*}_{l}$, can also be taken with respect to the posterior restricted to its region of attraction to shed light on which points are more uncertain (or stable) in their cluster allocation within the region of attraction.
@@
The contribution to the EVI for the meet of the particles is displayed in Figure 3(c) (bottom). By analyzing the patterns in this figure, we notice that the points with the largest EVIC are the ones at the border between the four large clusters (depicted in Figure 3(c) (top) and represented here with different shapes). Moreover, the points in the top-left component (labeled with 1) have a slightly larger EVIC than the points in the bottom-right component (labeled with 6), reflecting that their cluster allocation changes more across the MCMC samples, while the cluster assignment of the points in cluster 6 change less.